# Spin precession effects in the phasing formula of eccentric compact binary inspirals till the second post-Newtonian order

### v1

@@ II.2 Eccentricity evolution
$+\left(\frac{y_{0}}{y}\right)^{19/3}\biggl[\left\{-\frac{1243}{11904}-\frac{89\delta}{32}\kappa_{A}+\frac{21703\nu}{5952}+\kappa_{S}\!\left(-\frac{89}{32}+\frac{89\nu}{16}\right)\right\}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)^{2}$
$+\left\{-\frac{1243\delta}{5952}-\frac{89\delta}{16}\kappa_{S}+\kappa_{A}\!\left(-\frac{89}{16}+\frac{89\nu}{8}\right)\right\}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)$
$+\left\{-\frac{1243}{11904}-\frac{89\delta}{32}\kappa_{A}-\frac{19217\nu}{5952}+\kappa_{S}\!\left(-\frac{89}{32}+\frac{89\nu}{16}\right)\right\}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)^{2}$
$+\left\{-\frac{123}{3968}+\frac{89\delta}{96}\kappa_{A}+\kappa_{S}\!\left(\frac{89}{96}-\frac{89\nu}{48}\right)+\frac{123\nu}{1984}\right\}\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}$
$+\left\{-\frac{123\delta}{1984}+\frac{89\delta}{48}\kappa_{S}+\kappa_{A}\!\left(\frac{89}{48}-\frac{89\nu}{24}\right)\right\}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}$
@@
$+\left\{-\frac{123}{3968}+\frac{89\delta}{96}\kappa_{A}+\kappa_{S}\!\left(\frac{89}{96}-\frac{89\nu}{48}\right)+\frac{123\nu}{1984}\right\}\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr]$
@@
$+\left(\frac{y_{0}}{y}\right)^{31/3}\biggl[\left\{\frac{13}{114}+\frac{89\delta}{32}\kappa_{A}+\kappa_{S}\!\left(\frac{89}{32}-\frac{89\nu}{16}\right)-\frac{3343\nu}{912}\right\}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)^{2}$
$+\left\{\frac{13\delta}{57}+\frac{89\delta}{16}\kappa_{S}+\kappa_{A}\!\left(\frac{89}{16}-\frac{89\nu}{8}\right)\right\}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)$
$+\left\{\frac{13}{114}+\frac{89\delta}{32}\kappa_{A}+\kappa_{S}\!\left(\frac{89}{32}-\frac{89\nu}{16}\right)+\frac{2927\nu}{912}\right\}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)^{2}$
$+\left\{\frac{13}{608}-\frac{89\delta}{96}\kappa_{A}-\frac{13\nu}{304}+\kappa_{S}\!\left(-\frac{89}{96}+\frac{89\nu}{48}\right)\right\}\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}$
$+\left\{\frac{13\delta}{304}-\frac{89\delta}{48}\kappa_{S}+\kappa_{A}\!\left(-\frac{89}{48}+\frac{89\nu}{24}\right)\right\}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}$
@@
$+\left\{\frac{13}{608}-\frac{89\delta}{96}\kappa_{A}-\frac{13\nu}{304}+\kappa_{S}\!\left(-\frac{89}{96}+\frac{89\nu}{48}\right)\right\}\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr]$ (15d)
@@
The solution till $\mathcal{O}\left(e_{0}^{4}\right)$ is then obtained by dividing (9) by $e^{2}$ and then substituting the $e_{0}^{2}$ order solution in the RHS of the DE, taking a Taylor series in the PN parameter till the desired PN order, and then finally multiplying the entire RHS by $e^{2}$ to make a separable DE, which is then solved by separation of variables method. This process is then continued till $\mathcal{O}\left(e_{0}^{8}\right)$. The solution $e^{2}(y)$ till $\mathcal{O}\left(e_{0}^{8}\right)$ is given in the supplementary material.
@@ II.3 Spin precession effects in TaylorT2
Upon inverting and integrating Eq. 2, we find the coalescence time to have the following structure-
@@
where the coefficients were found to be as follows-
$t_{\rm NS,circ}=$ $1+y^{4}\biggl(\frac{617\nu^{2}}{72}+\frac{5429\nu}{504}+\frac{3058673}{508032}\biggr)-\frac{32\pi y^{3}}{5}+y^{2}\biggl(\frac{11\nu}{3}+\frac{743}{252}\biggr)\,$, (17a)
$t_{\rm NS,ecc}=$ $\frac{15}{43}\biggl(\frac{y_{0}}{y}\biggr)^{19/3}+y^{2}\biggl[\biggl(\frac{y_{0}}{y}\biggr)^{25/3}\biggl(\frac{14165}{14448}-\frac{985\nu}{516}\biggr)+\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggl(\frac{2015\nu}{148}-\frac{166175}{12432}\biggr)\biggr]+y^{3}\biggl[\frac{1885\pi}{1032}\biggl(\frac{y_{0}}{y}\biggr)^{28/3}$
$-\frac{8293\pi}{816}\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggr]+y^{4}\biggl[\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl(\frac{90775\nu^{2}}{18576}-\frac{331585\nu}{130032}-\frac{5966255}{43690752}\biggr)+\biggl(\frac{y_{0}}{y}\biggr)^{25/3}\biggl(-\frac{396955\nu^{2}}{5328}$
$+\frac{6232745\nu}{55944}-\frac{470773775}{12531456}\biggr)+\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggl(\frac{1176803\nu^{2}}{13392}-\frac{2363633\nu}{20832}+\frac{277964969}{31497984}\biggr)\biggr]\,$, (17b)
$t_{\rm SO,circ}=$ $\frac{226\delta}{15}\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}+\biggl(\frac{226}{15}-\frac{152\nu}{15}\biggr)\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\,$, (17c)
$t_{\rm SO,ecc}=$ $\biggl(\frac{y_{0}}{y}\biggr)^{28/3}\biggl[\biggl(\frac{275\nu}{387}-\frac{785}{774}\biggr)\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}-\frac{785\delta}{774}\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr]$
$+\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggl[\frac{2444\delta}{153}\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}+\biggl(\frac{2444}{153}-\frac{1319\nu}{306}\biggr)\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr]\,$, (17d)
$t_{\rm SS,circ}=$ $\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-15\delta\kappa_{A}+\frac{19369\nu}{992}+(30\nu-15)\kappa_{S}+\frac{223}{1984}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-15\delta\kappa_{A}-\frac{19815\nu}{992}$
$+(30\nu-15)\kappa_{S}+\frac{223}{1984}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[(60\nu-30)\kappa_{A}+\frac{223\delta}{992}-30\delta\kappa_{S}\biggr]$
$+\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl[\biggl(\frac{75}{21328}-\frac{75\nu}{10664}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}+\frac{75\delta}{10664}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)+\biggl(\frac{75\nu}{10664}$
$-\frac{75}{21328}\biggr)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}-\frac{75\delta}{10664}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}+\biggl(\frac{75}{21328}-\frac{75\nu}{10664}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}+\biggl(\frac{75\nu}{10664}$
$-\frac{75}{21328}\biggr)\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[5\delta\kappa_{A}+\frac{471\nu}{992}+(5-10\nu)\kappa_{S}-\frac{471}{1984}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[(10-20\nu)\kappa_{A}-\frac{471\delta}{992}+10\delta\kappa_{S}\biggr]+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[5\delta\kappa_{A}+\frac{471\nu}{992}+(5-10\nu)\kappa_{S}-\frac{471}{1984}\biggr]\,$, (17e)
$t_{\rm SS,ecc}=$ $\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[\frac{1335\delta\kappa_{A}}{1376}-\frac{16715\nu}{13072}+\biggl(\frac{1335}{1376}-\frac{1335\nu}{688}\biggr)\kappa_{S}+\frac{65}{1634}\biggr]+$
$\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[\biggl(\frac{1335}{688}-\frac{1335\nu}{344}\biggr)\kappa_{A}+\frac{65\delta}{817}+\frac{1335\delta\kappa_{S}}{688}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\chi_{S,0}\biggr)^{2}\biggl[\frac{1335\delta\kappa_{A}}{1376}+\frac{14635\nu}{13072}$
$+\biggl(\frac{1335}{1376}-\frac{1335\nu}{688}\biggr)\kappa_{S}+\frac{65}{1634}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[-\frac{445\delta\kappa_{A}}{1376}-\frac{195\nu}{13072}+\biggl(\frac{445\nu}{688}-\frac{445}{1376}\biggr)\kappa_{S}+\frac{195}{26144}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{445\nu}{344}-\frac{445}{688}\biggr)\kappa_{A}+\frac{195\delta}{13072}-\frac{445\delta\kappa_{S}}{688}\biggr]+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[-\frac{445\delta\kappa_{A}}{1376}-\frac{195\nu}{13072}+\biggl(\frac{445\nu}{688}$
$-\frac{445}{1376}\biggr)\kappa_{S}+\frac{195}{26144}\biggr]\biggr\}+\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{15447\delta\kappa_{A}}{992}+\frac{1244921\nu}{61504}+\biggl(\frac{15447\nu}{496}-\frac{15447}{992}\biggr)\kappa_{S}$
$+\frac{7603}{123008}\biggr]+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggl)\biggl[\biggl(\frac{15447\nu}{248}-\frac{15447}{496}\biggr)\kappa_{A}+\frac{7603\delta}{61504}-\frac{15447\delta\kappa_{S}}{496}\biggr]$
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{15447\delta\kappa_{A}}{992}-\frac{1260127\nu}{61504}+\biggl(\frac{15447\nu}{496}-\frac{15447}{992}\biggr)\kappa_{S}+\frac{7603}{123008}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{5149\delta\kappa_{A}}{992}$
$+\frac{32031\nu}{61504}+\biggl(\frac{5149}{992}-\frac{5149\nu}{496}\biggr)\kappa_{S}-\frac{32031}{123008}\biggr]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{5149}{496}-\frac{5149\nu}{248}\biggr)\kappa_{A}-\frac{32031\delta}{61504}+\frac{5149\delta\kappa_{S}}{496}\biggr]$
$+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{5149\delta\kappa_{A}}{992}+\frac{32031\nu}{61504}+\biggl(\frac{5149}{992}-\frac{5149\nu}{496}\biggr)\kappa_{S}-\frac{32031}{123008}\biggr]\biggr\}+\biggl(\frac{y_{0}}{y}\biggr)^{50/3}\biggl\{\biggl(\frac{37575}{4674304}$
$-\frac{37575\nu}{2337152}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}+\frac{37575\delta}{2337152}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)+\biggl(\frac{37575\nu}{2337152}-\frac{37575}{4674304}\biggr)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}$
$-\frac{37575\delta}{2337152}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}+\biggl(\frac{37575}{4674304}-\frac{37575\nu}{2337152}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}+\biggl(\frac{37575\nu}{2337152}-\frac{37575}{4674304}\biggr)\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr\}$ (17f)
The TaylorT2 phase was obtained to be the following
@@
where,
$\phi_{\rm NS,circ}=$ $1+y^{4}\biggl(\frac{3085\nu^{2}}{144}+\frac{27145\nu}{1008}+\frac{15293365}{1016064}\biggr)-10\pi y^{3}+y^{2}\biggl(\frac{55\nu}{12}+\frac{3715}{1008}\biggr)\,$, (19a)
$\phi_{\rm NS,ecc}=$ $-\frac{105}{272}\biggl(\frac{y_{0}}{y}\biggr)^{19/3}+y^{2}\biggl[\biggl(\frac{6895\nu}{3264}-\frac{14165}{13056}\biggr)\biggl(\frac{y_{0}}{y}\biggr)^{25/3}+\biggl(\frac{4155\nu}{896}-\frac{794635}{75264}\biggr)\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggr]$
$+y^{3}\biggl[-\frac{13195\pi}{6528}\biggl(\frac{y_{0}}{y}\biggr)^{28/3}-\frac{313}{960}\pi\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggr]+y^{4}\biggl[\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl(-\frac{635425\nu^{2}}{117504}+\frac{19505\nu}{6912}+\frac{5966255}{39481344}\biggr)$
$+\biggl(\frac{y_{0}}{y}\biggr)^{25/3}\biggl(-\frac{272845\nu^{2}}{10752}+\frac{23982055\nu}{338688}-\frac{2251200955}{75866112}\biggr)+\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggl(\frac{3141415\nu^{2}}{76032}-\frac{30486055\nu}{354816}$
$+\frac{459664885}{178827264}\biggr)\biggr]\,$, (19b)
$\phi_{\rm SO,circ}=$ $\frac{565\delta}{24}\,\,\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}+\biggl(\frac{565}{24}-\frac{95\nu}{6}\biggr)\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\,$, (19c)
$\phi_{\rm SO,ecc}=$ $\biggl(\frac{y_{0}}{y}\biggr)^{28/3}\biggl[\frac{5495\delta}{4896}\,\,\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}+\biggl(\frac{5495}{4896}-\frac{1925\nu}{2448}\biggr)\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr]$
$+\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggl[\frac{179\delta}{72}\,\,\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}+\biggl(\frac{1393\nu}{360}+\frac{179}{72}\biggr)\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr]\,$, (19d)

### v2

@@ II.2 Eccentricity evolution
$\quad+\left(\frac{y_{0}}{y}\right)^{19/3}\biggl[\left\{-\frac{1243}{11904}-\frac{89\delta}{32}\kappa_{A}+\frac{5149\nu}{2976}+\kappa_{S}\!\left(-\frac{89}{32}+\frac{89\nu}{16}\right)\right\}\left(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\right)^{2}$
$\quad+\left\{-\frac{1243\delta}{5952}-\frac{89\delta}{16}\kappa_{S}+\kappa_{A}\!\left(-\frac{89}{16}+\frac{89\nu}{8}\right)\right\}\left(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\right)\left(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\right)$
$\quad+\left\{-\frac{1243}{11904}-\frac{89\delta}{32}\kappa_{A}-\frac{21\nu}{16}+\kappa_{S}\!\left(-\frac{89}{32}+\frac{89\nu}{16}\right)\right\}\left(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\right)^{2}$
$\quad+\left\{-\frac{123}{3968}+\frac{89\delta}{96}\kappa_{A}+\kappa_{S}\!\left(\frac{89}{96}-\frac{89\nu}{48}\right)+\frac{5887\nu}{2976}\right\}\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{A,0}$
$\quad+\left\{-\frac{123\delta}{1984}+\frac{89\delta}{48}\kappa_{S}+\kappa_{A}\!\left(\frac{89}{48}-\frac{89\nu}{24}\right)\right\}\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{S,0}$
@@
$\quad+\left\{-\frac{123}{3968}+\frac{89\delta}{96}\kappa_{A}+\kappa_{S}\!\left(\frac{89}{96}-\frac{89\nu}{48}\right)-\frac{89\nu}{48}\right\}\bm{\chi}_{S,0}\!\cdot\!\bm{\chi}_{S,0}\biggr]$
@@
$\quad+\left(\frac{y_{0}}{y}\right)^{31/3}\biggl[\left\{\frac{13}{114}+\frac{89\delta}{32}\kappa_{A}+\kappa_{S}\!\left(\frac{89}{32}-\frac{89\nu}{16}\right)-\frac{1613\nu}{912}\right\}\left(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\right)^{2}$
$\quad+\left\{\frac{13\delta}{57}+\frac{89\delta}{16}\kappa_{S}+\kappa_{A}\!\left(\frac{89}{16}-\frac{89\nu}{8}\right)\right\}\left(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\right)\left(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\right)$
$\quad+\left\{\frac{13}{114}+\frac{89\delta}{32}\kappa_{A}+\kappa_{S}\!\left(\frac{89}{32}-\frac{89\nu}{16}\right)+\frac{21\nu}{16}\right\}\left(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\right)^{2}$
$\quad+\left\{\frac{13}{608}-\frac{89\delta}{96}\kappa_{A}-\frac{1769\nu}{912}+\kappa_{S}\!\left(-\frac{89}{96}+\frac{89\nu}{48}\right)\right\}\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{A,0}$
$\quad+\left\{\frac{13\delta}{304}-\frac{89\delta}{48}\kappa_{S}+\kappa_{A}\!\left(-\frac{89}{48}+\frac{89\nu}{24}\right)\right\}\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{S,0}$
@@
$\quad+\left\{\frac{13}{608}-\frac{89\delta}{96}\kappa_{A}+\frac{89\nu}{48}+\kappa_{S}\!\left(-\frac{89}{96}+\frac{89\nu}{48}\right)\right\}\bm{\chi}_{S,0}\!\cdot\!\bm{\chi}_{S,0}\biggr]$ (15b)
@@
The solution up to $\mathcal{O}\left(e_{0}^{4}\right)$ is then obtained by dividing (9) by $e^{2}$ and then substituting the $e_{0}^{2}$ order solution in the RHS of the DE, taking a Taylor series in the PN parameter up to the desired PN order, and then finally multiplying the entire RHS by $e^{2}$ to make a separable DE, which is then solved by separation of variables method. This process is then continued up to $\mathcal{O}\left(e_{0}^{8}\right)$. The solution $e^{2}(y)$ up to $\mathcal{O}\left(e_{0}^{8}\right)$ is given in [104].
@@ II.3 Spin precession effects in TaylorT2
Upon inverting and integrating Eq. 2, we find the coalescence time, to have the following structure-
@@
where $t_{\rm NS,circ}$ and $t_{\rm NS,ecc}$ are provided in [105], $t_{\rm SO,circ}$ in [106], and $t_{\rm SO,ecc}$ in [101]. Rest of the coefficients are as follows-
$t_{\rm SS,circ}$ $=\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\biggr)^{2}\biggl[-15\delta\kappa_{A}+\frac{4489\nu}{496}+(30\nu-15)\kappa_{S}+\frac{223}{1984}\biggr]$
$+\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\biggr)^{2}\biggl[-15\delta\kappa_{A}-\frac{19\nu}{2}+(30\nu-15)\kappa_{S}+\frac{223}{1984}\biggr]$
$+\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\biggr)\biggl[(60\nu-30)\kappa_{A}+\frac{223\delta}{992}-30\delta\kappa_{S}\biggr]$
$+\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{A,0}\biggl[5\delta\kappa_{A}+\frac{5431\nu}{496}+(5-10\nu)\kappa_{S}-\frac{471}{1984}\biggr]$
$+\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{S,0}\biggl[(10-20\nu)\kappa_{A}-\frac{471\delta}{992}+10\delta\kappa_{S}\biggr]$
$+\bm{\chi}_{S,0}\!\cdot\!\bm{\chi}_{S,0}\biggl[5\delta\kappa_{A}-10\nu+(5-10\nu)\kappa_{S}-\frac{471}{1984}\biggr]$
$+\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl[\biggl(\frac{75}{21328}-\frac{75\nu}{5332}\biggr)\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\biggr)^{2}+\frac{75\delta}{10664}\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\biggr)$
$\quad+\frac{75}{21328}\biggl(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\biggr)^{2}+\biggl(\frac{75\nu}{5332}-\frac{75}{21328}\biggr)\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{A,0}$
$\quad-\frac{75\delta}{10664}\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{S,0}-\frac{75}{21328}\bm{\chi}_{S,0}\!\cdot\!\bm{\chi}_{S,0}\biggr]\,$, (17a)
$t_{\rm SS,ecc}$ $=\biggl(\frac{y_{0}}{y}\biggr)^{31/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[\frac{1335\delta\kappa_{A}}{1376}-\frac{8065\nu}{13072}+\biggl(\frac{1335}{1376}-\frac{1335\nu}{688}\biggr)\kappa_{S}+\frac{65}{1634}\biggr]$
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[\biggl(\frac{1335}{688}-\frac{1335\nu}{344}\biggr)\kappa_{A}+\frac{65\delta}{817}+\frac{1335\delta\kappa_{S}}{688}\biggr]$
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[\frac{1335\delta\kappa_{A}}{1376}+\frac{315\nu}{688}+\biggl(\frac{1335}{1376}-\frac{1335\nu}{688}\biggr)\kappa_{S}+\frac{65}{1634}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[-\frac{445\delta\kappa_{A}}{1376}-\frac{8845\nu}{13072}+\biggl(\frac{445\nu}{688}-\frac{445}{1376}\biggr)\kappa_{S}+\frac{195}{26144}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{445\nu}{344}-\frac{445}{688}\biggr)\kappa_{A}+\frac{195\delta}{13072}-\frac{445\delta\kappa_{S}}{688}\biggr]$
$+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[-\frac{445\delta\kappa_{A}}{1376}+\frac{445\nu}{688}+\biggl(\frac{445\nu}{688}-\frac{445}{1376}\biggr)\kappa_{S}+\frac{195}{26144}\biggr]\biggr\}$
$+\biggl(\frac{y_{0}}{y}\biggr)^{19/3}\biggl\{\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}\biggl[-\frac{15447\delta\kappa_{A}}{992}+\frac{287207\nu}{30752}+\biggl(\frac{15447\nu}{496}-\frac{15447}{992}\biggr)\kappa_{S}+\frac{7603}{123008}\biggr]$
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)\biggl[\biggl(\frac{15447\nu}{248}-\frac{15447}{496}\biggr)\kappa_{A}+\frac{7603\delta}{61504}-\frac{15447\delta\kappa_{S}}{496}\biggr]$
$+\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}\biggl[-\frac{15447\delta\kappa_{A}}{992}-\frac{4755\nu}{496}+\biggl(\frac{15447\nu}{496}-\frac{15447}{992}\biggr)\kappa_{S}+\frac{7603}{123008}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\biggl[\frac{5149\delta\kappa_{A}}{992}+\frac{351269\nu}{30752}+\biggl(\frac{5149}{992}-\frac{5149\nu}{496}\biggr)\kappa_{S}-\frac{32031}{123008}\biggr]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\biggl[\biggl(\frac{5149}{496}-\frac{5149\nu}{248}\biggr)\kappa_{A}-\frac{32031\delta}{61504}+\frac{5149\delta\kappa_{S}}{496}\biggr]$
$+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggl[\frac{5149\delta\kappa_{A}}{992}-\frac{5149\nu}{496}+\biggl(\frac{5149}{992}-\frac{5149\nu}{496}\biggr)\kappa_{S}-\frac{32031}{123008}\biggr]\biggr\}$
$+\biggl(\frac{y_{0}}{y}\biggr)^{50/3}\biggl\{\biggl(\frac{37575}{4674304}-\frac{37575\nu}{1168576}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)^{2}+\frac{37575\delta}{2337152}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\biggr)\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)$
$+\biggl(-\frac{37575}{4674304}+\frac{37575\nu}{1168576}\biggr)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}-\frac{37575\delta}{2337152}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}$
$+\frac{37575}{4674304}\biggl(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\biggr)^{2}-\frac{37575}{4674304}\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr\}.$ (17b)
Note that these coefficients up to $\mathcal{O}(e_{0}^{8})$ are given in [104].
The TaylorT2 phasing $\phi$ follows an evolution equation as follows-
$\frac{d\langle\phi\rangle}{dy}=\frac{d\langle\phi\rangle}{dt}\frac{dt}{dy}=\left(1-e^{2}\right)^{3/2}{y^{3}\over M}\frac{dt}{dy}\,$, (18)
Eq. 18 can be integrated using $dt/dy$ from the inversion of Eq. 2 to obtain the TaylorT2 phasing as a function of $e_{0}$ and $y$. It takes the following form-
@@
where $\phi_{\rm NS,circ}$ and $\phi_{\rm NS,ecc}$ are provided in [105], $\phi_{\rm SO,circ}$ in [106], and $\phi_{\rm SO,ecc}$ in [101]. Rest of the coefficients are as follows-